OpenStudy (anonymous):

Integrate 2*(tan^-1(x))^2 From x= 0 to 1 need help

OpenStudy (anonymous):

@hartnn -can u help with this one :)

OpenStudy (anonymous):

\[\int\limits_{0}^{1} 2(\tan ^{-1}x)^2 \]

OpenStudy (anonymous):

is it so?

OpenStudy (anonymous):

ok so first can u express the equation into simpler means?

OpenStudy (anonymous):

for example taking out the 2?

OpenStudy (anonymous):

\[[\int\limits] \arctan x (dx) = x \arctan x - (1/2) \ln(1+x2) + C\]

OpenStudy (anonymous):

integrate it now

OpenStudy (kc_kennylau):

Is the question even correct lol School won't give you questions that hard

OpenStudy (amoodarya):

if u=arctan x then you have int (0 to pi/4) 2u^2 (1+tan^2 u )du

OpenStudy (amoodarya):

then answer is .490563

OpenStudy (anonymous):

sorry question is wrong there was a mistake in the book it should be tan^-1(x^2)

OpenStudy (anonymous):

@Loser66-thanx for the help the minor misatke has made the question diffiuclt to solve it but u still did it :)

OpenStudy (anonymous):

@shakesworde- question was wrong :)

OpenStudy (anonymous):

@kc-kennylau-lol yeah school don't give hard questions :)

OpenStudy (anonymous):

@kc_kennylau *